Missions   Polar   Operations




What's New:

2005/01/21: There was a GTM anomaly at 021:04:00:20, perhaps due to the present solar flare. The ADC was reset at 16:07:31. The flipper reset was completed at 17:27.

2005/01/04: There was a GTM anomaly at 004:03:05:35, resulting in garbled telemetry to several instruments. The MFE ADC was upset.

2004/04/07: The flip status of MFE has been corrected to the FLIP LEFT direction at 2028 UT.

A GTM anomaly occurred at 094/09:04:52. The GTM was restored when the anomaly was first observed during the contact at 094/17:40-19:10 (April 4). MFE was recovered during the same contact. However flipper became undetermined after the anomaly.

2004/03/22: The flipper status of the MFE has been corrected as of 2000-2005 UT today.

2003/04/10: Although MFE had been returning magnetic field data after the A-D reset executed on April 7, it was noted that the instrument's flip status had been undetermined since the anomaly, as opposed to the expected flip-left position. This suggests that the anomaly impacted more than the A-D converter, and an RTC for correcting the flipper status was then performed during 2118-2138 UT on this day. In the RTC the on-board offset values were also re-installed to ensure their proper values. An examination of the MFE data suggests that the on-board offsets were erased between 1933 UT of 2003/04/07 and 2138 UT of 2003/04/10. The MFE has been operating normally after the RTC.

2003/04/07: At 15:13 UT the MFE instrument started returning constant magnetic field counts of (-32447, 32528, 32318) in the three components. The housekeeping parameters show that an anomaly in A-D occurred at that time, but the cause of this anomaly has not been identified. The instrument's A-D state was reset at 19:33 UT on the same day in an NRT interval, and the MFE data became normal afterward.


Polar MFE continuous to function normally.

The sensors maintain in the flip-left orientation.

None of the cross-strapped redundancy has yet been used.

Data are obtained from the inboard sensor in its high range at radial distances below about 3.12 Re. Above this distance data are obtained from the outboard sensor. Outboard gain is switched to the low range at a distance of about 6.24 Re.

Prior Major Events





Visitor - Research - Library - Education - Missions - Data - Personnel - Resources - News - Help

For more information about Space Science Center contact: www@igpp.ucla.edu.

Last updated: April 15, 2003